OpenStudy (anonymous):

six times the reciprocal of a number equals 2 times the reciprocal of 6. find the number.

OpenStudy (tkhunny):

Write a clear and concise defintion. This will save you from wandering around. N = The number we seek. Now translate... "six times the reciprocal of a number " Our number is "N". It's reciprocal is "1/N" Six times that is "6/N" "two times the reciprocal of 6 " 2/6 = 1/3 6/N = 1/3 - Solve away!

OpenStudy (anonymous):

N=2?

OpenStudy (tkhunny):

Please solve the equation using proper steps and eliminating guessing.

OpenStudy (anonymous):

N=18

OpenStudy (anonymous):

Thank you!

OpenStudy (tkhunny):

That's better. Guessing can be helpful, but we must be able to abandon our guess when the mathematics suggests otherwise. Good work.